ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 6, 15 Niezgoda, Marek, Institute of Applied Mathematics, Agric. Univ. of Lublin, Poland An unified approach to some majorization inequalities Group majorization theory is used to give an unified approach to some classical and recently proved majorization inequalities involving eigenvalues or singular values of matrices. An analog of the inequalities is derived for a class of group induced cone orderings. Reference: M. Niezgoda, Group Majorization and Schur Type Inequalities, Linear Algebra and Its Application 268 (1998), 9-30.